OpenStudy (anonymous):

tan*2*2x x tan*2*3x x tan5x = tan*2*2x - tan*2*3x + tan 5x equation

OpenStudy (anonymous):

Ok Is it \[tan2 \times 2x \]

OpenStudy (anonymous):

I mean can I write it as \[tan4x \times tan6x \times tan5x\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):